Find the flux through a spherical Gaussian surface of radius a=1 m surrounding a charge of 8.85 pC

1
Expert's answer
2022-03-09T12:19:25-0500

The Gauss's law says


"\Phi=\oint EdA=\frac{q}{\epsilon_0}""\Phi=\frac{8.85*10^{-12}}{8.85*10^{-12}}=1.00\:\rm N\cdot m^2/C"